### v3

@@ 1 Introduction
Figure 1: Overview and comparison of our All-Seeing Model v2 with other MLLMs. To enhance relation comprehension ability while maintaining grounding, referring, and other general capabilities, we propose (1) a novel task, termed Relation Conversation (ReC), which unifies the formulation of text generation, object localization, and relation comprehension; (2) a high-quality dataset AS-V2, which consists of over 127K samples for ReC; (3) the All-Seeing Model v2 (ASMv2), which is capable of comprehending and grounding the relations between the objects in the image.
@@ 2.1 Vision-Language Models
Significant advancements have been made in the field of visual recognition and understanding in recent years. Models based on the image-text matching framework [66, 29, 17, 11] achieve powerful zero-shot performance on various downstream tasks, thereby initiating the era of open-world semantic recognition and understanding. Subsequent works [38, 88] further integrate this framework with language modeling tasks to support more generative tasks. The recent progress of Large Language Models [5, 61, 78] leads to the emergency of many LLM-based multimodal models [36, 107, 51, 82, 39, 95, 11, 77, 90], which aim to integrate the powerful understanding and reasoning ability of LLMs with multimodal perception and comprehension. Despite their powerful performance, these works are only capable of capturing the holistic visual information of the whole image. Some recent methods [8, 63, 99, 81, 33, 96, 67, 53, 62] begin to focus on location-aware understanding. However, due to the lack of appropriate modeling methods and training data for relation comprehension, these methods struggle to comprehend the inter-object relations within images accurately. To enhance relation comprehension ability while maintaining other general capabilities of MLLMs, we introduce a novel task, termed Relation Conversation, which unifies the formulation of text generation, object localization, and relation comprehension.
@@ 3.1 The All-Seeing Dataset v2
The prompt for GPT-4V comprises six components: (1) Task description, which explains the formulation of relation conversation. (2) Image to be annotated. (3) Caption annotations of this image, intended to enhance GPT-4V’s understanding of the scene. (4) Location annotations, which are the bounding boxes of the objects in the scene and guide GPT-4V in annotating the objects in the caption. (5) Relation annotations, which are presented as a list of (subject, predicate, object) triplets and help GPT-4V to annotate the predicate in the caption. (6) Seed examples, which are manually annotated to assist GPT-4V in comprehending the task description and formatting the output. Although the caption annotations are not necessary for GPT-4V to produce the desired relation conversation data, incorporating these details into the prompt significantly reduces the hallucinations in the generated data. An example of the prompt is presented in Appendix 0.B.
The generated data comprise three types, including: (1) Detailed description, which requires the model to generate a comprehensive description for an image. (2) Region captioning, which requires the model to generate a comprehensive description for a certain region within the image. (3) Conversation, which requires the model to respond to the user query in the multi-turn conversation. The question types include the relations between objects, the object types, counting the objects, object actions, object locations, and relative positions between objects. Each type of data is generated using different task descriptions and human-annotated seed examples. These tasks require the model to understand pointer instructions (e.g., utilizing bounding boxes as prompts) and link the objects and predicates mentioned in the generated response to the image regions. An example is shown in Fig. 3, with additional examples in Appendix 0.B.
@@ 3.2 Circular-based Relation Probing Evaluation
Figure 4: Data examples in the CRPE. The benchmark is formulated as single-choice questions and consists of four splits: Existence, Subject, Predicate, and Object. The Existence split is designed to evaluate the object recognition ability while the remaining splits focus on the evaluation of relation comprehension. A qualitative comparison in abnormal data between LLaVA-1.5 and ASMv2 is shown at the bottom.
@@
Additionally, to evaluate the dependency on language priors, we further include abnormal data in the Predicate split, which depict relations that are rare but reasonable in the real world. Specifically, we first select relation triplets with minimal $P(p|s,o)$ and then employ DALLE-3 [3] to generate corresponding images for these triplets. Considering that the generated images might not match the specified triplets exactly, we perform a manual filtering process for these triplet-image pairs to ensure data quality. After that, we generate the evaluation data using the method mentioned above. Fig. 4 shows abnormal examples at the bottom and more examples of these abnormal data are presented in Appendix 0.D.
@@ 4.1 Relation Conversation
Figure 5: Data formulation for Relation Conversation. Each object is marked with <ref></ref> and followed by a bounding box indicating its location while each predicate is marked with <pred></pred> and followed by two bounding boxes referring to its subjects and objects.
@@ 5.2 Results on Region-level Benchmarks
Our ASMv2 achieves state-of-the-art performance on the representative REC benchmarks [31, 58]. As shown in Tab. 2, our ASMv2 significantly outperforms current state-of-the-art MLLMs, including Qwen-VL [2] and Ferret [86]. Specifically, our model surpasses Ferret by an average of 1.85 points across all benchmarks.
@@
ASMv2 (ours) 87.8 88.8 78.4
*ASMv2 (ours) 88.4 89.9 79.4
@@ 5.3 Results on Open-ended Scene Graph Generation
In this section, we evaluate the Relation Conversation capability of our model through the Open-ended Scene Graph Generation task on the Panoptic Scene Graph (PSG) dataset [85], which is a widely-used benchmark for the scene graph generation. See Sec. 0.C.2 for the results on the Predicate Classification task.
@@
As shown in Tab. 5, our ASMv2 demonstrates state-of-the-art performance in the OpenSGG task. Specifically, our ASMv2 significantly outperforms TextPSG by 8.7 points in recall while generating a significantly fewer average number of tuples compared to it (9.2 vs. 100.0). Note that having more tuples generally implies an advantage in computing recall. When compared to traditional scene graph generation models, which generate scene graphs in a close-ended manner, ASMv2 also exhibits competitive performance. Despite generating fewer tuples, our model maintains a competitive recall of 14.2 and a mean recall of 10.3. Another factor negatively impacting the performance is that our ASMv2 generates scene graphs in an open-ended manner while recall is calculated in an exact-match manner. Therefore, the triplets (people, standing on, grass) and (person, standing on, grass) are considered mismatched even though they represent the same semantics. A more appropriate metric for the OpenSGG task will be left for future work.

### v4

@@ 1 Introduction
Figure 1: Overview and comparison of our All-Seeing Model v2 with other MLLMs. In this project, we propose (1) a novel task, termed Relation Conversation (ReC), which unifies the formulation of text generation, object localization, and relation comprehension; (2) a high-quality dataset AS-V2, which consists of over 127K samples for ReC; (3) the All-Seeing Model v2 (ASMv2), which is capable of comprehending and grounding the relations between the objects in the image.
@@ 2.1 Vision-Language Models
Significant advancements have been made in the field of visual recognition and understanding in recent years. Models based on the image-text matching framework [68, 29, 17, 11] achieve powerful zero-shot performance on various downstream tasks, thereby initiating the era of open-world semantic recognition and understanding. Subsequent works [39, 90] further integrate this framework with language modeling tasks to support more generative tasks. The recent progress of Large Language Models [5, 63, 80] leads to the emergency of many LLM-based multimodal models [37, 109, 53, 84, 40, 97, 11, 79, 92, 30, 49], which aim to integrate the powerful understanding and reasoning ability of LLMs with multimodal perception and comprehension. Despite their powerful performance, these works are only capable of capturing the holistic visual information of the whole image. Some recent methods [8, 65, 101, 83, 34, 98, 69, 55, 64] begin to focus on location-aware understanding. However, due to the lack of appropriate modeling methods and training data for relation comprehension, these methods struggle to comprehend the inter-object relations within images accurately. To enhance relation comprehension ability while maintaining other general capabilities of MLLMs, we introduce a novel task, termed Relation Conversation, which unifies the formulation of text generation, object localization, and relation comprehension.
@@ 3.1 The All-Seeing Dataset v2
The prompt for GPT-4V comprises six components: (1) Task description, which explains the formulation of relation conversation. (2) Image to be annotated. (3) Caption annotations of this image, intended to enhance GPT-4V’s understanding of the scene. (4) Location annotations, which are the bounding boxes of the objects in the scene and guide GPT-4V in annotating the objects in the caption. (5) Relation annotations, which are presented as a list of (subject, predicate, object) triplets and help GPT-4V to annotate the predicate in the caption. (6) Seed examples, which are manually annotated to assist GPT-4V in comprehending the task description and formatting the output. Although the caption annotations are not necessary for GPT-4V to produce the desired relation conversation data, incorporating these details into the prompt significantly reduces the hallucinations in the generated data. An example of the prompt is presented in Appendix B.
The generated data comprise three types, including: (1) Detailed description, which requires the model to generate a comprehensive description for an image. (2) Region captioning, which requires the model to generate a comprehensive description for a certain region within the image. (3) Conversation, which requires the model to respond to the user query in the multi-turn conversation. The question types include the relations between objects, the object types, counting the objects, object actions, object locations, and relative positions between objects. Each type of data is generated using different task descriptions and human-annotated seed examples. These tasks require the model to understand pointer instructions (e.g., utilizing bounding boxes as prompts) and link the objects and predicates mentioned in the generated response to the image regions. An example is shown in Fig. 3, with more examples in Appendix B.
@@ 3.2 Circular-based Relation Probing Evaluation
Figure 4: Data examples in the CRPE. The benchmark consists of four splits: Existence, Subject, Predicate, and Object. A qualitative comparison in abnormal data between LLaVA-1.5 and ASMv2 is shown at the bottom.
@@
Additionally, to evaluate the dependency on language priors, we further include abnormal data in the Predicate split, which depict relations that are rare but reasonable in the real world. Specifically, we first select relation triplets with minimal $P(p|s,o)$ and then employ DALLE-3 [3] to generate corresponding images for these triplets. Considering that the generated images might not match the specified triplets exactly, we perform a manual filtering process for these triplet-image pairs to ensure data quality. After that, we generate the evaluation data using the method mentioned above. Fig. 4 shows abnormal examples at the bottom and more examples are presented in Appendix D.
@@ 4.1 Relation Conversation
Figure 5: Data formulation for Relation Conversation. Each object is marked with <ref></ref> and followed by a box denoting its location while each predicate is marked with <pred></pred> and followed by two boxes referring to its subjects and objects.
@@ 5.2 Results on Region-level Benchmarks
Our ASMv2 achieves state-of-the-art performance on the representative REC benchmarks [32, 60]. As shown in Tab. 2, our ASMv2 significantly outperforms current state-of-the-art MLLMs, including Qwen-VL [2] and Ferret [88].
@@
ASMv2-13B (ours) 87.8 88.8 78.4
*ASMv2-13B (ours) 88.4 89.9 79.4
@@ 5.3 Results on Open-ended Scene Graph Generation
In this section, we evaluate the Relation Conversation capability of our model through the Open-ended Scene Graph Generation task on the Panoptic Scene Graph (PSG) dataset [87], which is a widely-used benchmark for the scene graph generation. See Appendix C.2 for the results on the Predicate Classification task.
@@
As shown in Tab. 5, our ASMv2 demonstrates state-of-the-art performance in the OpenSGG task. Specifically, our ASMv2 significantly outperforms TextPSG by 8.7 points in recall while generating a significantly fewer average number of tuples compared to it (9.2 vs. 100.0). Note that having more tuples generally implies an advantage in computing recall. When compared to traditional scene graph generation models, which generate scene graphs in a close-ended manner, ASMv2 also exhibits competitive performance. Despite generating fewer tuples, our model maintains a competitive recall of 14.2 and a mean recall of 10.3. Another factor negatively impacting the performance is that our ASMv2 generates scene graphs in an open-ended manner while recall is calculated in an exact-match manner. Therefore, the triplets (people, standing on, grass) and (person, standing on, grass) are considered mismatched even though they represent the same semantics. A more appropriate metric for this task will be left for future work.
@@ 6 Conclusion
Acknowledgments
The work is supported by the National Key R&D Program of China (NO. 2022ZD0161300), and the National Natural Science Foundation of China (Grant No. 62376134).